# Examining Inequality in Park Quality for Promoting Health Across 35 Global Cities

### v1

@@ Abstract
Urban parks provide significant health benefits by offering spaces and facilities for various recreational and leisure activities. However, the capacity of specific park spaces (e.g., children’s playgrounds, lakes) and elements (e.g., benches, sports equipment) to foster health remains underexamined. Traditional studies have focused on parks’ size, greenery, and accessibility, often overlooking their ability to facilitate specific health-promoting activities. To address this gap, we propose a taxonomy consisting of six categories of health-promoting activities in parks: physical, mind-body, nature appreciation, environmental, social, and cultural. We estimate the capacity of parks in 35 global cities to promote health by establishing a lexicon linking park spaces and elements with specific health-promoting activities from our taxonomy. Using this lexicon, we collected data on elements and spaces in all parks in 35 cities from OpenStreetMap. Our analysis covers 23,477 parks with a total of 827,038 elements and spaces. By first comparing similarly sized parks across cities, we found that North American parks offer more spaces for physical activities, while European parks focus more on nature appreciation. Second, by scoring parks based on both elements and spaces, we investigated the variability in their health-promoting potential. We found the most uniform provision across parks for physical activities and the highest disparities regarding social activities. Additionally, parks offering a variety of activities are usually located in city centers, while offerings diminish in parks towards the suburbs. Lastly, we identified significant inequalities in park standards across cities, regardless of their continental location: Tokyo and Paris offer the most uniform park standards, while Copenhagen and Rio de Janeiro exhibit the most pronounced disparities. Our study provides insights for making urban parks more equitable, engaging, and health-promoting.
@@
1 Introduction
As the world’s population continues to gravitate towards urban areas[1, 2], cities are faced with the immense task of creating and maintaining green spaces to foster public health[3, 4, 5]. Urban green spaces are shown to enhance various health aspects, including physical, mental, and social[6, 7, 8, 9, 10, 11, 12, 13]. Among specific benefits, having access to parks lowers rates of morbidity[14], reduces social isolation[15], decreases stress levels[16], and contributes to overall well-being[17]. Health benefits of parks are especially relevant to socioeconomically deprived [6, 14] and elderly populations[18, 19, 20].
Park developers seeking cost-effective ways to promote health [21, 22, 23, 24] have placed strong emphasis on physical activity[25, 26, 27, 28]. This is not surprising, given that physical inactivity has been identified as a major health issue in Western countries[29] leading to record levels of obesity in the United States[30]. There are also worrying levels of child obesity in Chinese mega cities [31]. However, beyond physical activities, even mere exposure to the natural environment and its appreciation has been shown beneficial for health[32]. For instance, research by Olszewska-Guizzo et al.[33] highlights how the visual characteristics of green spaces can elicit positive emotional responses and brain activities akin to stress reduction, yet such aesthetic considerations have seldom been prioritized in park design[34]. Social interactions, too, are critical for overall health and well-being[35], leading in some countries to the adoption of the practice of social prescribing[36]. Parks can complement other urban environments in providing engaging social as well as cultural experiences[37], with literature indicating potential strategies for park management to facilitate these experiences[38].
Despite these various activities through which parks can promote health, previous work linking urban greenery with health has mainly considered parks as mere green boxes, e.g., looking at the size and type of green space in them or at their distance from concentrated human habitation[39, 40, 41], and did not dive deeper into the specific activities afforded by those parks. Example factors identified to drive park visitations, in general, include the area, number of food establishments, and the presence of natural vegetation and fields[42]. Other works have focused on the surroundings of parks, such as accessibility and the socioeconomic environment, factors which were then used to estimate the park usage through location-based social network check-in data[43]. Overall, the attributes to understand park usage in prior work are quite broad and disregard the amount of and specific facilities [44, 42, 43, 45]. In summary, prior research on the connection between parks and health largely overlooked the range of potential health-promoting activities these spaces can offer. Additionally, studies focusing on park usage patterns have only examined a limited selection of facilities without adequately connecting these facilities to the usage and health-promoting activities.
To address this gap and provide a detailed understanding of park quality with respect to various activities that people can do in them, we collected data on park elements and spaces in parks across 35 cities worldwide. The raw counts reveal design priorities for park activities on different continents, while the park health scores indicate a trend where parks closer to city centers offer a wider range of facilities. Additionally, we identified the successes and failures of various cities in providing high-quality parks, highlighting those that offer more uniformly distributed park facilities compared to others that show significant disparities in facility provision.
@@ 2 Results
We first present the lexicon of health-promoting park facilities that we operationalized with OpenStreetMap (OSM). We then focus on 35 large cities listed in Table 3, selected to cover various parts of the world and to have good Internet connectivity and OSM data availability, allowing for robust analyses. Afterward, we present the results of scoring parks in those cities based on their facilities, highlighting the globally outstanding parks we uncover for each activity category. We conclude with an inequality analysis of parks both within and across cities.
2.1 Lexicon of Health-Promoting Park Facilities
We aimed at identifying health benefits associated with specific activities performed in parks and measuring park facilities’ support for these activities. To achieve this, we implemented a three-step method (detailed in Methods Section 4.3): 1) identifying activities in parks, 2) confirming health benefits of these activities, and 3) gathering park facility data and connecting it with health-promoting activities.
Step 1: Identifying Activities in Parks.
We convened an expert panel to identify common park activities, which they organized into six main categories, constituting a taxonomy of activities in parks (described in Methods Section 4.3, Expert Panel):
Physical activities: moving your body and sports, like walking, hiking, biking, swimming, and group fitness classes;
Mind-body activities: practices that combine movement, breathing, and meditation for relaxation and well-being. Examples include yoga, meditation, and tai chi;
@@ 2.1 Lexicon of Health-Promoting Park Facilities
Nature-appreciation activities: enjoying nature, with activities like bird watching, camping, picnicking, and nature journaling;
Environmental activities: gardening and park conservation, such as planting trees and flowers or participating in conservation efforts;
Social activities: coming together with others, like attending outdoor events, playing sports, and volunteering; and
Cultural activities: celebrating community diversity and heritage through activities like cultural festivals, art exhibits, music performances, storytelling, and workshops.
The categories of this taxonomy sometimes overlap. For instance, gardening falls under both environmental and nature-appreciation, and team sports, though physical, also have a social aspect. The expert panel agreed that, while there are overlaps, the taxonomy is accurate and valuable, comprehensively covering a broad range of park activities.
Activity Category Health Aspect Specific Health Benefit
Physical Cognitive health dementia prevention[46]
General health longevity[47, 19]
Social health various[96, 60], social cohesion[97, 55, 98, 58]
General health lower morbidity[101]
General health lower morbidity[101]
Social health access to healthy produce[132], social cohesion[128, 132]
General health longevity[19]
Cultural Cognitive health dementia prevention[46]
General health various[148]
Physical health increase of physical activity[149]
Well-being quality of life[46, 147, 149]
Mind-body Mental health stress reduction[112], anxiety reduction[56], depression prevention[59, 56]
Physical health increase of physical activity[90]
Well-being quality of life[150]
Step 2: Confirming Health Benefits of these Activities.
A literature review (detailed in Methods Section 4.3, Scoping Review) was done to link these park activities to health benefits. After identifying 114 research studies as relevant, we annotated each study with the activity category and the specific health benefit this study evidenced. We found that engaging in activities within urban parks provides a range of health benefits, impacting physical, mental, and social well-being (Table 1).
Most research has focused on the benefits of physical activity in parks. Out of 79 studies on the health benefits of exercising, 46 underscored positive outcomes like weight loss[62, 63, 64], cardiovascular improvements[91, 93], metabolic activity[70, 71]. Additionally, these activities demonstrated positive effects on mental health (16 articles), well-being (7 articles), and social health (6 articles).
Finally, cultural and mind-body activities are relatively under-researched regarding their health benefits in the context of urban parks. Cultural activities often fell outside the scope of our review, which required a connection to urban parks, while more general cultural activities were studied. However, their health benefits are likely underreported given the presence of cultural facilities in parks (e.g., historic monuments or arts venues). Yoga and other mind-body exercises have not been studied as much in the context of parks, highlighting a potential gap in the scientific literature that warrants future exploration.
Step 3: Gathering Park Facility Data and Connecting it with Health-promoting Activities.
Activity Category Elements Spaces
Using OSM, a collaborative mapping platform, we collected all parks in the 35 cities and curated all map objects within the outlines of these parks. This established a large collection of park facility data, which are either designated spaces, such as a forest or pond, or elements such as a bench or an individual tree. Each element and space is described by tags, which we used to assign them to one of the activity categories. To scale this approach, we employed a large language model (LLM)-based classifier, which we validated using an expert-derived gold standard dataset (Section 4.4, Using Large Language Models for Annotation). We display the top 10 most frequent tags for each activity category in Table 2 and publish all 1441 entries of the lexicon in the replication package. As there were no OSM tags that supported mind-body activities (see Section 4.4, Operationalization of the Taxonomy), this category was not considered in the rest of this study, but could be considered by future work if behavioral data on this activity category would be available.
The collected OSM tags allow to categorize different elements and spaces in parks, giving us a detailed insight into their potential use and linked health benefits. This new way of examining urban parks allowed us to compare parks within and across different cities based on their health-promoting potential.
2.2 Park Health Scores
To quantify a park’s health-promoting potential, we measured how well it can support each of the health-promoting activities from our taxonomy. To do so, we scored each park based on the presence of the gathered facilities (OSM elements and spaces) that were associated with each specific activity relative to the park’s size. In the end, we combined the counts of elements and areas covered by spaces into one overall score for each health activity category (Section 4.5).
The main outcome of our study is the scoring of 23,477 parks across 35 cities in five continents based on their potential for health-promoting activities. Each park’s score captures its health-promoting offering relative to an average park of the same size in the city. As an illustrative example, Figure 1 shows scores for nature appreciation and physical activities in London parks. The park health scores of all 35 cities are visualized in an interactive online map at https://social-dynamics.net/healthy-parks/visualization/.
Figure 1: Visualization of the park health scores in London, UK. On the left, scores for nature appreciation are shown, and on the right, scores for physical activities. The diverging color palette uses green for high nature appreciation scores, red for high physical health scores, white for average park offerings, and black for low scores. The web interface allows browsing individual park scores in each of the study’s cities: https://social-dynamics.net/healthy-parks/visualization/.
Quantitative Validation Through Geolocated Flickr Images
Table 3: Correlation between OSM park health scores and Flickr activity scores. The values showcase the connection of available facilities and photographed usage of parks. The results are based on all parks with at least 250 Flickr images within their boundaries.
@@ 2.2 Park Health Scores
As our primary validation of the park health activity scores, we verified whether activities seen in millions of individual Flickr images in parks closely matched the potential activities derived from OSM data. We used a global dataset of geotagged photos from Flickr, providing an alternative way to evaluate park usage [42]. We measured Flickr park activity using both user-generated tags and automatic tags generated from computer vision algorithms that label objects within the images. By semantically mapping the Flickr tags using sentence embeddings to OSM tags, we transferred our activity taxonomy to the domain of Flickr tags (Methods Section 4.7). Matching the tags allowed us to characterize parks based on the visual content within the photos and calculate activity scores similar to those computed from OSM.
To determine how well the OSM and Flickr scores aligned across different activity categories, we used the Pearson Correlation Coefficent (PCC) (Table 3). Overall, the correlation coefficients suggest that the amount of images that people share on Flickr shows the greatest agreement with the facilities present in parks on OSM for cultural attractions ($\mu=0.53$), with a PCC ranging from $0.30$ in Amsterdam up to $0.75$ in Philadelphia; followed by social activities ($\mu=0.39$), ranging from $0.15$ in Buenos Aires up to $0.72$ in Chicago; and environmental activities ($\mu=0.39$), ranging from $0.01$ in Copenhagen, up to $0.93$ in Stockholm. The correlation is lowest for nature-appreciation ($\mu=0.33$), ranging from $-0.04$ in Madrid up to $0.72$ in Perth, and physical activities ($\mu=0.30$), ranging from $-0.04$ in Buenos Aires up to $0.47$ in Seattle. These results are influenced by the socio-cultural nature of Flickr [151] and motivations for tagging photos[152]. The average correlation across 35 cities worldwide demonstrates that the OSM scores are in line with the photographed uses of our parks. We found an average PCC of $0.39$, with a fairly low standard deviation of $0.07$. The strongest links between OSM and Flickr activity scores were seen in Washington DC ($\mu=0.52$) and Perth ($\mu=0.50$). Conversely, the weakest correlations, at $0.30$, were found in Amsterdam, Hong Kong, and Vienna. Despite the fact that we translated tags from local languages into English, a pattern emerged: cities where English is predominantly spoken tended to have stronger correlations between the two data sources. We suspect this might be attributable to a range of factors, including that the OSM and Flickr data, primarily lacking from non-English-speaking countries, could lead to both, limited usage of two platforms and reduced data quality in those regions.
Qualitative Validation of Globally Outstanding Parks
Table 4: Top five parks globally for each health-promoting activity category as well as an overall score. We have linked each park to its Wikipedia entry or municipal webpage.
Activity Category City Park Score
Physical Sydney Centennial Park 6.90
Paris Bois de Boulogne 6.83
Taipeh Da’an Forest Park ({CJK}UTF8min大安森林公園) 6.47
Tokyo Wakasu Seaside Park ({CJK}UTF8min若洲海浜公園) 6.08
Rio De Janeiro Aterro do Flamengo 5.71
Nature-appreciation Madrid Dehesa de la Villa 6.59
Tokyo Ueno Park ({CJK}UTF8min上野恩賜公園) 6.35
Taipeh Guandu Nature Park ({CJK}UTF8min關渡自然公園) 6.23
Rio De Janeiro Quinta da Boa Vista 6.14
Montreal Parc Angrignon 5.64
Environmental Seattle Washington Park Arboretum 4.95
Tokyo Kita-no-maru Park ({CJK}UTF8min北の丸公園) 4.19
Christchurch Hagley Park North 3.84
New York Bronx Park 3.70
Madrid Retiro Park 3.36
Social Moscow Timiryazevskiy Park (Тимирязевский парк) 4.82
Hong Kong Inspiration Lake Recreation Centre {CJK}UTF8min(迪欣湖活動中心) 4.79
Toronto Toronto Island Park 4.32
Madrid Finca Vista Alegre 4.18
Tokyo Ueno Park {CJK}UTF8min(上野恩賜公園) 4.08
Cultural Tokyo Ueno Park {CJK}UTF8min(上野恩賜公園) 8.20
Seattle Seattle Center 4.61
Washington DC National Mall 4.24
Seoul Gyeongbokgung Palace Park ({CJK}UTF8mj경복궁) 4.19
St Petersburg Garden of Saint Petersburg State Forestry University (парк Лесотехнической академии) 4.15
Overall Paris Bois de Boulogne 3.82
Tokyo {CJK} UTF8min Ueno Park (上野恩賜公園) 3.74
Washington DC National Mall 3.74
Christchurch Hagley Park North 3.62
Paris Bois de Vincennes 3.57
As a second way to validate our methods, we identified the most outstanding parks in terms of the health benefits they offer. We found that urban parks that score high in a specific category, such as social or cultural, are internationally renowned for facilities in that specific category, as evidenced by their Wikipedia pages. To do so, we picked out and evaluated the parks with the highest scores in each of the five categories (Section 2.2). The full list of the top parks for each city and category is shown in Table 8 in the Appendix. These parks are renowned globally as natural, cultural, sports, and social landmarks in their cities, famous for the facilities they offer. To initially validate the associations of those parks with their activity categories, we ascertained whether the parks’ Wikipedia pages would discuss the facilities supporting activities in those categories.
In the physical category, top parks serve as sports centers. Australia’s Centennial Parklands and Paris’s Bois de Boulogne offer a variety of sports facilities, such as fields for tennis, soccer, polo, and race courses. Rio’s Flamengo Park is known for beach sports, and Tokyo’s Wakasu Seaside Park, a 2020 Olympics venue, provides ample space for sports, such as golf.
Nature-appreciation parks, like Madrid’s La Dehesa de la Villa and Taipei’s Guandu Nature Park, are valued for their lush vegetation and wildlife. Tokyo’s Ueno Park and Montreal’s Parc Angrignon are celebrated for their dense greenery and biodiversity.
For environmental activities, parks with extensive botanical gardens stand out. The Washington Park Arboretum in Seattle and Kita-no-maru Park in Tokyo focus on plant preservation and include unique gardens. Hagley Park in Christchurch and Bronx Park in New York feature premier botanical gardens, while Madrid’s Retiro Park offers a green oasis with community gardening spaces.
In the social sphere, parks with diverse entertainment and social spaces excel. Hong Kong’s Inspiration Lake offers leisure activities, while Toronto Island Park features amenities for social gatherings. Ueno Park in Tokyo is also a popular spot for its cafés and social spaces, showcasing parks as essential places for community and leisure.
In the cultural category of Section 2.2, we spotlighted parks rich in history and culture. Tokyo’s Ueno Park, spotlighted for the third time, is filled with historic sites like the Kan’ei-ji Temple, along with the Tokyo National Museum. Seattle Center is a cultural hub with museums and the Seattle Opera. The National Mall in Washington DC is packed with museums and monuments. Seoul’s Gyeongbokgung Palace Park hosts the ancient royal palace and museums. The Garden at the Saint Petersburg State Forestry University is known for its educational and historic significance.
Finally, in the overall category, which represents the mean value of the five other categories, we identified parks that stand out in their respective cities for their comprehensive offerings across all dimensions. These are typically large parks of outstanding significance, receiving the utmost attention from their municipal maintainers (e.g., Bois de Boulogne in Paris and, once more, Ueno Park in Tokyo).
To further confirm our method’s effectiveness in identifying leading parks for various categories, we conducted online searches using queries structured as: “[city name] parks for [activity category] activities” (e.g., “Tokyo parks for cultural activities”). Consistently, articles about the parks our method pinpointed appeared within the top ten search results for each category (e.g., “Ueno Park: A Seasonal Guide to Tokyo’s Cultural Haven”).
Average Park Offerings
By comparing the offerings in parks across different continents, we reveal varying priorities in urban planning and park design for supporting health-promoting activities. Figure 2 shows the expected number of elements and the size of spaces for individual activity categories in a fictional, “statistically average” 8-hectare park for cities in the different continents. We chose this size as it represents a medium-sized neighborhood park and is also the mean park size in our study. The values are computed using the parameters of the linear regression models that we used to score the parks for each city (see details in Section 4.5, Transforming Counts into Health Scores). Since we fit individual linear models for each city, our health scores for individual parks in different cities are incomparable, however, the statistical model we used allowed for the comparison of the baseline parks falling on the “average park line” across different cities (cf. Figure 7).
Figure 2 reveals that nature-appreciation and physical spaces are most common. When it comes to elements, nature-appreciation is again the most frequent category, followed by social and environmental categories. Cultural elements and spaces are the least frequent ones. In the Appendix, Figure 8, we showcase the same data subdivided by cities.
@@
Figure 2: Total area of park spaces (horizontal axis) versus the count of elements (vertical axis) in an 8-hectare park for each activity, broken down by continents. Nature-appreciation and physical activity spaces are the most common, with elements for nature-appreciation also outnumbering other activity categories. Cultural elements and spaces are the least frequent.

### v2

# Health-promoting Potential of Parks in 35 Cities Worldwide

@@ Abstract
Urban parks are important for public health, but the role of specific spaces, such as playgrounds or lakes, and elements, such as benches or sports equipment, in supporting well-being is not well understood. Based on expert input and a review of the literature, we defined six types of health-related activities: physical, mindfulness, nature appreciation, environmental, social, and cultural. We built a lexicon that links each activity to specific elements and spaces within parks present in OpenStreetMap. Using this data, we scored 23,477 parks across 35 cities worldwide based on their ability to support these activities. We found clear patterns: parks in North America focus more on physical activity, while those in Europe offer more chances to enjoy nature. Parks near city centers support health-promoting activities better than those farther out. Suburban parks in many cities lack the spaces and equipment needed for nature-based, social, and cultural activities. We also found large gaps in park quality between cities. Tokyo and Paris provide more equal access, while Copenhagen and Rio de Janeiro show sharp contrasts. These results can help cities create fairer parks that better support public health.
@@
1 Main
As the world’s population continues to gravitate towards urban areas, cities are faced with the immense task of creating and maintaining green spaces to foster public health[1, 2]. Urban parks are especially beneficial to vulnerable population groups such as socioeconomically deprived[3] and older people[4, 5]. Yet the provision of amenities and facilities that support health-promoting activities has not been systematically analyzed at scale.
Urban green spaces can support health in five main ways. First, they help people stay active, which can reduce obesity and heart diseases[6, 7]. Second, they give people a calm space to rest, reduce stress, and improve focus [8, 9, 6, 10]. Third, they bring people together and support social ties [11]. Fourth, they clean the air, block noise, and cool the city [6, 12]. Last, some green areas expose people to biodiverse forms of life, which may improve the immune system [13, 14].
Park planners have often focused on practical ways to boost public health, such as adding sports fields, paths, or playgrounds [15, 16, 17, 18]. Despite groving evidence of positive effects, spaces that allowed a diverse range of people to relax or enjoy beauty were often neglected in the past[19]. Parks also have the power to build stronger communities[20], and many planners now recognize that parks offer clear environmental gains such as cleaner air and lower noise levels[12]. During the COVID-19 pandemic, cities saw how vital parks are for public well-being. This led to stronger calls for fair and flexible park design[21].
Most past studies have treated parks as simple green zones, based on size or distance from homes [22, 23]. Fewer have studied how park layout and equipment affect how people use them. For instance, SOPARC is a tool that tracks how people move through parks and what spaces they use [24]. Other tools look at what makes parks good for activity, such as access, safety, or looks [25]. Still others focus on blue spaces such as rivers or lakes [26]. These tools work well but are time-consuming to implement and hard to scale, and most focus only on physical activity [27, 28]. We still lack a clear understanding of the net benefits of parks, especially when weighing their positive aspects, such as providing spaces for recreation and supporting urban health[29, 3], with potential downsides, including reinforcing disparities [30, 31, 32] or contributing to gentrification[33, 34].
For many cities, there is no easy way to list what parks contain or link these parts to health uses. In this study, we score parks in 35 cities based on how well they support six types of health-related activity. We then ask how these scores relate to fair access to well equipped parks. Our work answers three questions. First, which health-related activities do different park spaces support? Second, how can we measure parks around the world based on this support? Third, where are the biggest gaps in what parks provide?
@@ 2 Results
We began by creating a lexicon that linked park spaces and elements to health-related activities using OpenStreetMap (OSM) data. We then scored parks in 35 major cities, which we selected for their geographic diversity and reliable OSM coverage. After validating these scores through statistical comparison and expert review, we analyzed differences in park offerings within and between cities.
2.1 Lexicon of Health-Related Park Features
Step 1: Defining Activities in Parks
An expert panel of three researchers with expertise in urban computing, Earth observation, and computational social science, identified common park activities and organized them into six categories:
Physical activities: movement and sport such as walking, biking, swimming, and group exercise;
Mindfulness activities: practices like yoga, meditation, and tai chi;
@@ 2.1 Lexicon of Health-Related Park Features
Nature appreciation: observing and enjoying the natural environment, including bird watching and picnicking;
Environmental activities: community involvement such as gardening and conservation;
Social activities: gatherings and group events such as festivals and volunteering;
Cultural activities: heritage and arts programs, including performances and exhibitions.
While our 6 categories occasionally overlapped, they captured park activities at a more useful granular level than the typical recreational/physical division. Some overlap is inevitable since certain activities like football (with social and physical aspects) are inherently multi-faceted.
Step 2: Linking Facilities to Activities
We used OSM to collect elements and spaces located within parks in the 35 cities. These included defined areas like forests and ponds as well as features such as benches and sports courts. Each item was described with a tag, which we used to assign the item to an activity category. To scale this process, we trained a classifier based on an large language model (LLM) and validated it with an expert-coded dataset. The full lexicon included $1,441$ OSM tags. The ten most frequent tags for each category are listed in Table 3 in Appendix. No OSM tags matched the mindfulness category, so we did not include it in our scoring. Future research may incorporate behavioral data to address this gap.
Figure 1: Scores for nature appreciation (left) and physical activities (right) in London parks. Some parks score low on nature appreciation but high on physical activity (A), while imperial-era parks such as Green Park and St. James’s Park, close to the royal palace score high on nature appreciation but low on physical activity (B). Queen Elizabeth Olympic Park, purpose-built for the 2012 Games, scores high across all dimensions (C).
Map data from OpenStreetMap https://www.openstreetmap.org/copyright
2.2 Scoring Parks for Health Support
We scored each park based on the number of activity-related elements and spaces it contained, normalized by park area. Each park received a score for every activity category. Figure 1 presents an example of these scores for London. In total, we scored 23,477 parks across 35 cities on five continents (Table 1). Each score reflects how well a park supports a specific activity compared to other parks of similar size in the same city.
Validation with Flickr Data
Table 1: Our activity scores correlate with the prevalence of photographed activities from Flickr images. Pearson Correlation Coefficients (PCC) between OSM scores and Flickr activity scores for parks with at least 250 photos.
@@ 2.2 Scoring Parks for Health Support
We validated our park scores using geotagged Flickr images taken within park boundaries. These images included both user-generated and machine-generated tags. We used sentence embeddings to match these tags to our activity taxonomy, allowing us to estimate activity scores for each park based on photo content.
We compared these Flickr-based activity scores with our OSM-derived scores using the Pearson Correlation Coefficent (PCC). Across the 35 cities, the average correlation was $0.39$ with a standard deviation of $0.07$, indicating moderate but consistent agreement between the two sources. The highest agreement appeared in the cultural ($\mu=0.53$), social ($\mu=0.39$), and environmental ($\mu=0.39$) categories. These types of activities often involve distinctive and photogenic features such as festivals, artworks, or gardens, which are likely to be photographed and tagged.
In contrast, the physical ($\mu=0.30$) and nature-appreciation ($\mu=0.33$) categories showed weaker correlations. We believe this may be because users often photograph people or scenery without tagging specific sports or nature-related elements, making it harder to detect those activities in the data.
At the city level, Washington, DC ($\mu=0.52$) and Perth ($\mu=0.50$) had the strongest correlations. These cities are in English-speaking countries where Flickr and OSM usage is more common, which likely led to higher data quality and tag relevance. Cities with lower correlations, including Amsterdam, Hong Kong, and Vienna (each $\mu\cong 0.3$), may have had fewer geotagged images, lower tag accuracy, or less alignment with our English-based taxonomy.
These findings support the use of OSM data to assess park infrastructure for health-promoting activities, especially in categories and regions where online content about park use is widely available.
Validation with Wikipedia
We also validated our activity scores using Wikipedia pages. We identified the top-scoring parks globally for each health-related activity based on our scoring system. For each park, we examined the Wikipedia page to confirm whether the park was known for the corresponding activity type.
In the physical category, we found that parks such as Centennial Parklands in Australia and Bois de Boulogne in Paris were known for extensive sports infrastructure, including fields for tennis, soccer, and polo. For nature appreciation, top parks like La Dehesa de la Villa in Madrid and Guandu Nature Park in Taipei stood out in their cities with large green areas and wildlife. In the environmental category, Washington Park Arboretum in Seattle and Kita-no-maru Park in Tokyo included botanical gardens and conservation efforts. Social activity leaders like Inspiration Lake in Hong Kong and Toronto Island Park in Canada were cited for hosting gatherings and public amenities. In the cultural category, Ueno Park in Tokyo featured museums and historic landmarks, while Seattle Center included concert halls and art venues.
To further verify our assignments, we conducted structured search queries, such as “Tokyo parks for cultural activities” and found that the top-ranked parks by our method consistently appeared in online recommendations and guides. A complete list of these parks, organized by city and activity category, is available in Table 4.
@@
Figure 2: Top: Area of park spaces (horizontal axis) and count of elements (vertical axis) in a typical 8-hectare park by activity and city. Nature and physical activity spaces are the most frequent. Nature-related elements are also the most common across cities. Cultural spaces and elements appear least often. Axes are scaled for clarity.